(1) Table 30

Country Gini Coefficient  
Sweden 0.24
France 0.275
Italy 0.319
United Kingdom 0.334
Portugal 0.395
Latvia 0.377

Refer to Table 30. Which country has the most unequal income distribution?

Latvia

Portugal

France

Sweden

(2) Table 20

Quantity of Number of Baseballs Labor Per Day
0 0
1 40
2 78
3 113
4 145
5 175

Refer to Table 20. This table describes the number of baseballs a manufacturing firm can produce per day with different quantities of labor. Each baseball sells for $5 in a competitive market. Suppose that the firm pays its workers $172 per day. To maximize its profit, how many workers should this firm hire?

1 worker

2 workers

3 workers

4 workers

5 workers

(3)Table 20

Quantity of Number of Baseballs Labor Per Day
0 0
1 20
2 48
3 72
4 92
5 108

Refer to Table 20. This table describes the number of baseballs a manufacturer can produce per day with different quantities of labor. Each baseball sells for $5.50 in a competitive market. What is the marginal revenue product of the fourth worker?

$20.

$88

$110

$132

$506

(4)

Table 28

Harold and Maude own a dance studio where they and their employees teach ballroom dancing. Their company is a competitive, profit-maximizing firm. Harold and Maude’s total production is detailed in the table below.

Labor
(# of workers) Output (in number of students per day)
0 0
1 90
2 170
3 220
4 250
5 270

Refer to Table 28. If Harold and Maude offer a special sale for one day where dance lessons are $5 each, and if they pay their workers $140 per day for that day, what is the maximum number of workers that they will hire for that day?

1

2

3

4

5

(5)

Table 21

Consider the following daily production data for MadeFromScratch, Inc. MadeFromScratch sells cupcakes for $2.40 each and pays the workers a wage of $186 per day.

Labor
(number of workers) Quantity (cupcakes per day) Wage (per day)
0 0 $186
1 97 $186
2 186 $186
3 272 $186
4 351 $186
5 426 $186

Refer to Table 21. To maximize the profit, how many workers should MadeFromScratch hire?

2 workers

3 workers

4 workers

5 workers
Answer #1

Question 1) Higher the value of gini coefficient, higher will be the inequality. So, Portugal has the higherest gini coefficient. SoC, option 2 is the correct answer.

Question 2) Marginal revenue product from employing the first labour is = 40*5 = $200

Marginal revenue product from employing the second labour is = 38*5 = $190

Marginal revenue product from employing the third labour is = (113-78)*5 = $175

Marginal revenue product from employing the fourth labour is = (145-113)*5 = $160

So, we can see that the marginal revenue product of fourth labour is less than the marginal cost (wage) of employing a labour, whereas the marginal revenue product of third labour is more than the wage. So, 3 workers should be employed. So option 3 is the correct answer.

Question 3) Marginal revenue product of fourth labour =

= Marginal product of 4th labour* price

= (92-72)*5.50

= $110.

So, option 3 is the correct answer.

Question 4) Marginal revenue product from employing the first labour is = 90*5 = $450

Marginal revenue product from employing the second labour is = (170-90)*5 = $400

Marginal revenue product from employing the third labour is = (220-170)*5 = $250

Marginal revenue product from employing the fourth labour is = (250-220)*5 = $150

Marginal revenue product from employing the fifth labour is = (270-250)*5 = $100.

So, we can see that the marginal revenue product of fifth labour is less than the marginal cost (wage) of employing a labour, whereas the marginal revenue product of fourth labour is more than the wage. So, 4 workers should be employed. So option 4 is the correct answer.

Question 5)

Marginal revenue product from employing the first labour is = 97*$2.4 = $232.8

Marginal revenue product from employing the second labour is = (186-97)*2.4 = $213.6

Marginal revenue product from employing the third labour is = (272-186)*2.4= $206.4

Marginal revenue product from employing the fourth labour is = (351-272)*2.4= $189.6

Marginal revenue product from employing the fifth labour is = (426-351)*2.4 = $180.

So, we can see that the marginal revenue product of fifth labour is less than the marginal cost (wage) of employing a labour, whereas the marginal revenue product of fourth labour is more than the wage. So, 4 workers should be employed. So option 3 is the correct answer.

IF YOU FIND THESE ANSWERS HELPFUL, PLEASE UP VOTE.